OpenStudy (anonymous):

Rachel went hiking for 15 minutes. Express this time in hour and minutes A)1hr 5min B)1hr 50min C)2hr 5min D)2hr 30min

OpenStudy (anonymous):

yeah same as @dmezzullo says @babbypower1o1

OpenStudy (anonymous):

15 or 150? if 150 then itds D

OpenStudy (hba):

I guess it is 150 min,Right @babbypower1o1

OpenStudy (aravindg):

I have 2 apples . The number if apples with me is a)20 b)4 c)3 d)100

Parth (parthkohli):

@AravindG 4.

OpenStudy (aravindg):

your question is similiar to the above one .can you check up again?

Parth (parthkohli):

When you cut out both apples in half :-D

OpenStudy (aravindg):

@ParthKohli gosh you solved it !

OpenStudy (aravindg):

@babbypower1o1 can you check the question again?

Parth (parthkohli):

Similarly, when you multiply by 10, you get the answer...

Parth (parthkohli):

@tayluvsya "Duh" is not allowed in mathematics. ;-)

OpenStudy (anonymous):

so many comments being typed o_o

OpenStudy (hba):

We know that there are 60 min in 1 hour. So, Divide 150/60=

OpenStudy (aravindg):

:) lets wait fo the right question.

Parth (parthkohli):

Q: Prove that \(a^2 +b^2 = c^2\). A: Duh. What?!

OpenStudy (anonymous):

thank you guys i was really needing to get that qustion done ty

OpenStudy (anonymous):

np

OpenStudy (hba):

You're welcome bro :)

OpenStudy (anonymous):

parth must have a long answer o_o

Parth (parthkohli):

Just for the sake of complicating it,\[\left\lfloor\dfrac{n}{60}\right\rfloor + n~\operatorname{modulo} 60\]Is how you convert \(n\) minutes to hours where \(+\) is the concatenation function. T_T